# DPER: Diffusion Prior Driven Neural Representation for Limited Angle and Sparse View CT Reconstruction

### v1

@@ 4.2 Comparison with SOTA Methods for LACT Task
We compare the proposed DPER model with the eight baselines on the AAPM and LIDC datasets for the LACT reconstructions with scanning angles of [0, 90]°, [0, 120]°, and [0, 150]° scanning ranges. Table 1 shows the quantitative results. On the AAPM dataset, our DPER, DiffusionMBIR and MCG produce the top-three performances and are significantly superior to the other baselines, benefiting from the generative prior provided by the diffusion model. However, on the LIDC dataset, we observe that DiffusionMBIR and MCG, trained on the AAPM dataset, experience severe performance degradation due to the out-of-distribution (OOD) problem. In some cases, their results were even worse than naive INR-based SCOPE, for example, 35.67 dB compared to 38.36 dB in PSNR for the [0, 150]° range. In contrast, the performance of our method remains stable over two different test datasets. Compared with DiffusionMBIR and MCG, DPER achieves superior results for the demanding OOD LIDC dataset by utilizing the INR method, which guarantees dependable data fidelity with the CT measurement data while maintaining the prior local image consistency.
Fig. 2 demonstrates the qualitative results on two representative samples (#0 and #1) from the AAPM and LIDC datasets. Upon visual inspection, it is evident that the model-based FBP and three INR-based methods (NePR, CoIL, and SCOPE) fail to yield satisfactory results on either dataset. This limitation is primarily attributed to the significant data gap inherent in the LACT problem. Although DCAR and DiffusionMBIR produce satisfactory results on the AAPM dataset, their reconstructions on the LIDC dataset exhibit noticeable distortions, primarily attributed to the out-of-distribution (OOD) problem. Conversely, the CT images generated by our DPER model across both datasets demonstrate high image quality in terms of both local details and global structural integrity.
@@ 4.3 Comparison with SOTA Methods for Ultra-SVCT Task
Additionally, we conduct a comparative evaluation of our DPER model against eight baseline methods on the AAPM and LIDC datasets for ultra-SVCT reconstructions with 10, 20, and 30 input views. We present the quantitative results in Table 2. The three INR-based methods hardly produce pleasing results for the ultra-SVCT task. Supervised FBPConvNet and DDNet perform well on the AAPM dataset while suffering from severe performance drops on the out-of-distribution LIDC dataset. DPER, DiffusionMBIR and MCG all obtain excellent and robust results on two datasets.
@@
Fig. 3 shows the qualitative result on two representative samples (#133 and #1) of AAPM and LIDC datasets. The results from the FBP algorithm suffer severe streaking artifacts, while INR-based methods (NeRP, CoIL, and SCOPE) produce overly smooth artifacts results. INR-based methods tend to overfit the incomplete measurement and lack data distribution prior information, resulting in less detailed tissue reconstruction compared to the supervised method FBPConvNet. FBPConvNet yields relatively acceptable results. However, the structure precision remains somewhat unclear. In the ultra-SVCT task, DPER and DiffusionMBIR obtain the best and second-best performance (Except for 10 views, DPER gets second-best PSNR and best SSIM). Compared to DiffusionMBIR, our DPER achieves comparable performance on the AAPM dataset, while on out-of-distribution LIDC data, we slightly outperform in data consistency and performance. The findings from LIDC show that the three diffusion-based methods produce favorable reconstruction results, even when dealing with out-of-distribution data in the ultra-SVCT task.
4.4 Clinical Dataset Results
@@
Figure 4: Comparison of LACT task (a. [0, 90]° and b. [0, 120]°) and SVCT task (c. 20 views) on test samples (#90, #178 and #213) of the clinical COVID-19 CT dataset.
@@
Table 3: Quantitative results (PSNR / SSIM) of different methods on the clinical COVID-19 CT dataset for LACT task (90° and 120°) and SVCT task (20 views). The Best performances are highlighted in bold.
Method LACT SVCT
[0, 90]° [0, 120]° 20 views
FBP 12.01 / 0.2370 14.48 / 0.2788 11.51 / 0.1492
DiffusionMBIR [35] 26.01 / 0.8815 28.54 / 0.9192 30.96 / 0.9119
DPER 31.63 / 0.9548 33.07 / 0.9678 31.92 / 0.9286
@@
To conduct a further comprehensive assessment on the effectiveness of our method, we undertaken experimental validation on the clinical COVID-19 CT dataset. We compared the proposed DPER with the current state-of-the-art method DiffusionMBIR on LACT task (with 90° and 120° scanning, respectively) and SVCT task (with 20 views projection).
@@
Fig. 4 shows the qualitative results on some samples (#90, #178 and #213) of the clinical COVID-19 CT dataset. Confronted with this challenging OOD scenario (different acquisition protocols, anomalous structures caused by diseases), the proposed DPER consistently demonstrates commendable generalization performance. As depicted in Fig. 4(a), for the 90° LACT task, our DPER significantly outperforms the SOTA method DiffusionMBIR, both in quantitative metrics and visual representation. Specifically, in areas highlighted by orange circles and blue arrows, the reconstruction results from DiffusionMBIR either omit or distort the original bone structure. In contrast, the results obtained from DPER closely align with the ground truth. This is further corroborated by the 120° LACT task results, as illustrated in Fig. 4(b). In situations involving anomalous structures attributable to COVID-19, our DPER method demonstrated a notable capability to recover the original content of these structures to the greatest extent possible. In contrast, the DiffusionMBIR method exhibited blurring and inaccuracies in reconstructing these anomalous areas. In summary, DPER can still hold outstanding performance in the more complex out-of-distribution clinical dataset without any fine-tuning on pre-trained diffusion prior.
@@ 4.5 Effectiveness of Refinement Interval
In the proposed framework, the hyperparameter refinement interval RI balances the contribution of the data fidelity sub-problem and the distribution prior sub-problem in Eq. (15). To evaluate its impact, we set four distinct RI values (25, 50, 100, and 200) on the AAPM dataset for the LACT task with the scanning range of [0, 90]°. Fig. 5 shows the model performance and time cost curve varying with the refinement interval and the qualitative comparisons. It is clear that as the refinement interval increases, the performance of the model gradually decreases, but the running time of the model increases considerably. As the refinement interval increases, the influence of the distribution prior regularization becomes more pronounced in the solution of the overall inverse problem, while the impact of the data fidelity subproblem diminishes. Thereby, the reconstructed image may be divergent from GT. The visualization results are consistent with the quantitative analysis. In short, refinement intervals provide us with a balanced choice between performance and time consumption. It is worth noting that when the $\texttt{RI}\leqslant 50$, further reduction of the refinement interval results in a very limited performance gain (e.g., only 0.14 improvement in PSNR) but a multiplicative computational cost. Therefore, in this paper we set $\texttt{RI}=50$.
@@ 4.6 Comparison of MBIR and INR as Data-fidelity Solver
Figure 6: Qualitative and quantitative results of MBIR (ADMM-TV) and INR with/without score-based function(DiffusionMBIR is ADMM-TV with score-based function, DPER is INR with score-based function) on one test sample(# 8)) of the AAPM dataset for the LACT task of [0, 90]° scanning range.
@@ 4.7 Effectiveness of Tweedie Denoising
To fully illustrate the benefits of Tweedie’s strategy for $\mathbf{x}_{0}$ estimation, we conducted ablation experiments to compare the difference between using $\mathbf{x}_{0}$ (i.e., $\left\|\mathbf{A}\hat{\mathbf{x}}_{0|t}-\mathbf{y}\right\|_{2}^{2}$) and $\mathbf{x}_{t}$ (i.e., $\left\|\mathbf{A}{\mathbf{x}}_{t}-\mathbf{y}\right\|_{2}^{2}$ for solving the data-fidelity subproblem. Fig. 7 displays the qualitative results of these two schemes on the AAPM dataset of the LACT task of [0, 90]° scanning range and SVCT task of 20 views. In the SVCT task, the reconstruction results of DPER without Tweedie denoising exhibit significant blur (as shown in the zoomed-in area of Fig. 7(a)). These results are characterized by indistinct boundaries between tissues and the presence of noticeable noise, which compromises the clarity and distinctness of the reconstructed images. In contrast, the reconstruction of DPER with Tweedie denoising is much clearer. The 90° LACT task shows a similar phenomenon; for example, in Fig. 7(b), the blue arrow indicates that Tweedie denoising provides more precise information. DPER with Tweedie outperforms DPER without Tweedie in both LACT and SVCT tasks. Recently, a lot of work has been demonstrating that using Tweedie denoising helps improve the performance of solving inverse problems via diffusion model [34, 47]. Indeed, using estimated $\mathbf{x}_{0}$ for data consistency operations (i.e., $\left\|\mathbf{A}\hat{\mathbf{x}}_{0|t}-\mathbf{y}\right\|_{2}^{2}$)) has become a widely used and recognized paradigm in the field of solving inverse problems via diffusion model [34, 47, 46, 42].
@@ 4.8 Effectiveness of Alternating Optimization Strategy
To evaluate the effectiveness of alternative optimization, we compared two naive strategies (as shown in Fig. 8) that directly integrate the INR and diffusion model [52, 53] without alternative optimization of the proposed framework. Strategy A: INR first represents the implicit representation field by fitting the incomplete measurement directly. Then, the pre-trained diffusion model performs conditional sampling to refine the represented field to produce the reconstruction result. Strategy B: First, the pre-trained diffusion model performs conditional sampling based on the FBP result of the incomplete measurement. Then, the generated result is refined by INR (followed by the Implicit Neural Representation Refinement process of DPER) to produce the reconstruction result. Conditional sampling is performed by incorporating the input image and the according level noise.
@@ 4.9.2 Influence of Noises
During the process of CT acquisition, the presence of noise is an inevitable phenomenon, attributable to a myriad of factors. Utilizing the limited angle sinogram $\mathbf{y}$, encompassing 90 views from the AAPM dataset, we simulate the measurement domain noise. This is accomplished through the employment of a statistical Poisson model, as delineated in the following:
@@
$\mathbf{Y(\theta,\rho)}$ represents the transmitted X-ray photon intensity, where $b$ denotes the incident X-ray photon intensity, and $r$ symbolizes the average of the background events and read-out noise variance. To simulate different Signal-to-Noise Ratio (SNR) levels, we assign the value of $r$ as 10, while $b$ is set at $1.3\times 10^{4}$, $4\times 10^{4}$, and $4\times 10^{5}$, corresponding to approximate SNR levels of 32dB, 37 dB, and 43dB, respectively. The resulting noisy sinograms, denoted as $\mathbf{y}_{\theta}^{\prime}$, are computed using the negative logarithm transformation: $\mathbf{y}_{\theta}^{\prime}=-\ln(\mathbf{Y}(\theta,\rho)/b)$. In our study, we evaluate the performance of our proposed DPER method against three baseline methodologies: FBP, SCOPE (as the SOTA in INR-based methods), and DiffusionMBIR (as the SOTA in diffusion-based methods) for CT image recovery from these noisy sinograms.
Figure 10: Qualitative and quantitative results of different methods to compare the measurement with the noise level of SNR $\approx$ 34 dB and without Noise on one test sample (#18) of the AAPM dataset for the LACT task of [0, 90]° scanning range.
Fig. 10 and Table 4 display the quantitative and qualitative results of the impact of different noise levels on our framework and other baselines for observing the robustness. There are some observations: 1) The results indicate that noise affects the model’s performance as it leads to a more ill-posed LACT inverse imaging problem. It is evident that as the noise level increases, the performance of all models deteriorates. However, the reduction ratio of self-supervised models shows better results than Prior-based methods, which is expected (The effect of diffusion prior is limited by noise measurement). 2) Compared to the FBP result with noise, the DPER result is significantly cleaner and maintains the data fidelity, which represents that the DPER framework can adapt to a single instance with different noise levels. 3) Compared to the SOTA DiffusionMBIR, our framework still maintains almost +2dB PSNR and +0.03 SSIM proportion of lead in all noise settings.
@@
Table 4: Quantitative results (PSNR / SSIM) of different methods on the AAPM dataset test samples for LACT task of [0,90]° scanning range with three levels of noise (SNR = 32 dB, 37dB, 43 dB). The Best performances are highlighted in bold.
Method 32 dB 37 dB 43 dB
FBP 11.43 / 0.2018 11.47 0.2366 11.51 / 0.2974
SCOPE[20] 22.75 / 0.6829 23.08 0.7409 23.32 / 0.8124
DiffusionMBIR[35] 31.88 / 0.9234 33.07 0.9390 34.15 / 0.9503
DPER (Ours) 33.36 / 0.9459 34.52 0.9602 36.41 / 0.9685

### v2

@@ 4.2 Comparison with SOTA Methods for LACT Task
We compare the proposed DPER model with the eight baselines on the AAPM and LIDC datasets for the LACT reconstructions with scanning angles of [0, 90]°, [0, 120]°, and [0, 150]° scanning ranges. Table 1 shows the quantitative results. On the AAPM dataset, our DPER, DiffusionMBIR and MCG produce the top-three performances and are significantly superior to the other baselines, benefiting from the generative prior provided by the diffusion model. However, on the LIDC dataset, we observe that DiffusionMBIR and MCG, trained on the AAPM dataset, experience severe performance degradation due to the OOD problem. In some cases, their results were even worse than naive INR-based SCOPE, for example, 35.67 dB compared to 38.36 dB in PSNR for the [0, 150]° range. In contrast, the performance of our method remains stable over two different test datasets. Compared with DiffusionMBIR and MCG, DPER achieves superior results for the demanding OOD LIDC dataset by utilizing the INR method, which guarantees dependable data fidelity with the CT measurement data while maintaining the prior local image consistency.
Fig. 2 demonstrates the qualitative results on two representative samples (#0 and #1) from the AAPM and LIDC datasets. Upon visual inspection, it is evident that the model-based FBP and three INR-based methods (NePR, CoIL, and SCOPE) fail to yield satisfactory results on either dataset. This limitation is primarily attributed to the significant data gap inherent in the LACT problem. Although DCAR and DiffusionMBIR produce satisfactory results on the AAPM dataset, their reconstructions on the LIDC dataset exhibit noticeable distortions, primarily attributed to the OOD problem. Conversely, the CT images generated by our DPER model across both datasets demonstrate high image quality in terms of both local details and global structural integrity.
@@ 4.3 Comparison with SOTA Methods for Ultra-SVCT Task
We conduct a comparative evaluation of our DPER model against eight baseline methods on the AAPM and LIDC datasets for ultra-SVCT reconstructions with 10, 20, and 30 input views. We present the quantitative results in Table 2. The three INR-based methods hardly produce pleasing results for the ultra-SVCT task. Supervised FBPConvNet and DDNet perform well on the AAPM dataset while suffering from severe performance drops on the OOD LIDC dataset. DPER, DiffusionMBIR and MCG all obtain excellent and robust results on two datasets.
@@
Fig. 3 shows the qualitative result on two representative samples (#133 and #1) of AAPM and LIDC datasets. The results from the FBP algorithm suffer severe streaking artifacts, while INR-based methods (NeRP, CoIL, and SCOPE) produce overly smooth artifacts results. INR-based methods tend to overfit the incomplete measurement and lack data distribution prior information, resulting in less detailed tissue reconstruction compared to the supervised method FBPConvNet. FBPConvNet yields relatively acceptable results. However, the structure precision remains somewhat unclear. In the ultra-SVCT task, DPER and DiffusionMBIR obtain the best and second-best performance. Compared to DiffusionMBIR, our DPER method is superior in both quantitative metrics and qualitative visualization, exhibiting less noise and more accurate structures. The findings from LIDC show that the three diffusion-based methods produce favorable reconstruction results, even when dealing with OOD data in the ultra-SVCT task.
4.4 Comparison on Clinical COVID-19 Dataset
@@
Figure 4: Comparison of LACT task ([0, 90]° and [0, 120]°) and SVCT task (20 views) on test samples (#90, #178 and #213) of the clinical COVID-19 CT dataset.
@@
Table 3: Quantitative results (PSNR / SSIM) of different methods on the clinical COVID-19 dataset and real LDCT-PD projection dataset for LACT task (90° and 120°) and SVCT task (20 views and 30 views). The Best performances are highlighted in bold.
Dataset Method LACT SVCT
[0, 90]° [0, 120]° 20 views 30 views
COVID-19 FBP [67] 12.01 / 0.2370 14.48 / 0.2788 11.51 / 0.1492 21.85 / 0.2700
MCG [45] 27.59 / 0.9153 29.08 / 0.9207  30.37 / 0.9082 32.14 / 0.9228
DiffusionMBIR [46] 26.73 / 0.8879 29.54 / 0.9292 30.96 / 0.9119 32.93 / 0.9314
DPER (Ours) 31.63 / 0.9548 33.07 / 0.9678 31.92 / 0.9286 34.75 / 0.9466
LDCT-PD FBP [67] 13.08 / 0.2814 16.09 / 0.3467 16.19 / 0.0979 18.55 / 0.1438
MCG [45] 23.61 / 0.8515 27.86 / 0.8916 29.08 / 0.8972 31.10 / 0.9038
DiffusionMBIR [46] 21.46 / 0.8482 25.52 / 0.8947 29.75 / 0.8937 31.60 / 0.9097
DPER (Ours) 28.21 / 0.9139 32.60 / 0.9439 30.84 / 0.9123 32.80 / 0.9226
@@
To conduct a more comprehensive assessment of the effectiveness of our method, we carried out experimental validation using the clinical COVID-19 CT dataset. We compared the proposed DPER with the current SOTA method MCG and DiffusionMBIR on LACT task (with 90° and 120° scanning, respectively) and SVCT task (with 20 views and 30 views projection, respectively). We present the quantitative results in Table 3.
@@
Fig. 4 shows the qualitative results on some samples (#90, #178 and #213) of the clinical COVID-19 CT dataset. Confronted with this challenging OOD scenario (different acquisition protocols, anomalous structures caused by diseases), the proposed DPER consistently demonstrates commendable generalization performance. As depicted in the first row in Fig. 4, for the 90° LACT task, our DPER significantly outperforms the SOTA methods DiffusionMBIR and MCG, both in quantitative metrics and visual representation. Specifically, in areas highlighted by pink circles and green arrows, the reconstruction results from DiffusionMBIR either omit or distort the original bone structure. In contrast, the results obtained from DPER closely align with the ground truth. This is further corroborated by the 120° LACT task results, as illustrated in Fig. 4. In cases presenting anomalous structures associated with COVID-19, our DPER method exhibited superior capability in reconstructing the original content of these structures with high fidelity. In contrast, the DiffusionMBIR method exhibited blurring and inaccuracies in reconstructing these anomalous areas. In summary, DPER can still hold outstanding performance in the more complex OOD clinical dataset without any fine-tuning on pre-trained diffusion prior.
4.5 Comparison on Real Projection Dataset
Figure 5: Comparison of LACT task ([0, 120]°) and SVCT task (30 views) on representative samples of the real projection clinical LDCT-PD dataset.
The proposed DPER is also evaluated on the real projections data from the LDCT-and-Projection-data dataset to more closely match challenging real-world clinical scenarios. Raw projection data makes the reconstruction task challenging due to non-eliminable noise in the acquisition process, complex acquisition protocols and even possible geometric errors. We compared the proposed DPER with two SOTA methods, MCG and DiffusionMBIR, on the same settings as the COVID-19 dataset.
Quantitative and qualitative results are displayed in Table 3. and Fig. 5, respectively. It is clear that the FBP results for both the LACT and SVCT tasks have difficulty in partitioning structure and detail due to severe artifacts. For the 120° LACT task, scrutinizing the specific ROIs in Fig. 5, we find that DiffusionMBIR is unable to restore the correct shape of the bone. Additionally, its reconstruction results show a severe defect in the lower left. This may be due to its departure from the prior manifold of the data, caused by the continuous forced application of data consistency operations. MCG obtains slightly better results with Manifold Constraints [45], but demonstrates poor data fidelity. In particular, the green arrows mark the wrong bone and muscle structures. In contrast, the proposed DPER is highly consistent with GT both in terms of prior and measurement manifold. Notably, a close inspection of the reconstruction results reveals that both DiffusionMBIR and MCG exhibit cartoon-like features that blur fine structures. This blurring is due to the disparity between the pre-trained score function and the target distribution of the reference image. The proposed DPER, on the other hand, can fully utilize the respective strengths of the data prior and measurement data to produce accurate reconstructions with clear details. The 30-views SVCT task shows similar findings. In the regions indicated by the pink circles and green arrows, both DiffusionMBIR and MCG introduced blurred or erroneous structures as well as significant noise, while DPER was closer to the reference image.
@@ 4.6 Effectiveness of Refinement Interval
In the proposed framework, the hyperparameter refinement interval RI balances the contribution of the data fidelity sub-problem and the distribution prior sub-problem in Eq. (22). To evaluate its impact, we set four distinct RI values (25, 50, 100, and 200) on the AAPM dataset for the LACT task with the scanning range of [0, 90]°. Fig. 6 shows the model performance and time cost curve varying with the refinement interval and the qualitative comparisons. It is clear that as the refinement interval increases, the performance of the model gradually decreases, but the running time of the model increases considerably. As the refinement interval increases, the influence of the distribution prior regularization becomes more pronounced in the solution of the overall inverse problem, while the impact of the data fidelity subproblem diminishes. Consequently, the reconstructed image may deviate from the ground truth. The visualization results are consistent with the quantitative analysis. In short, refinement intervals provide us with a balanced choice between performance and time consumption. It is worth noting that when the $\texttt{RI}\leqslant 50$, further reduction of the refinement interval results in a very limited performance gain (e.g., only 0.14 improvement in PSNR) but a multiplicative computational cost. Therefore, in this paper we set $\texttt{RI}=50$.
@@ 4.7 Comparison of MBIR and INR as Data-fidelity Solver
Figure 7: Qualitative and quantitative results of MBIR (ADMM-TV) and INR with and without score-based prior (DiffusionMBIR is ADMM-TV with score-based prior, and DPER is INR with score-based prior) on one test sample(# 8)) of the AAPM dataset for the LACT task of [0, 90]° scanning range.
@@ 4.8 Effectiveness of Tweedie Denoising
To fully illustrate the benefits of Tweedie’s strategy for $\mathbf{x}_{0}$ estimation, we conducted ablation experiments to compare the difference between using $\mathbf{x}_{0}$ (i.e., $\left\|\mathbf{A}\hat{\mathbf{x}}_{0|t}-\mathbf{y}\right\|_{2}^{2}$) and $\mathbf{x}_{t}$ (i.e., $\left\|\mathbf{A}{\mathbf{x}}_{t}-\mathbf{y}\right\|_{2}^{2}$ for solving the data-fidelity subproblem. Fig. 8 displays the qualitative results of these two schemes on the AAPM dataset of the LACT task of [0, 90]° scanning range and SVCT task of 20 views. In the SVCT task, the reconstruction results of DPER without Tweedie denoising exhibit significant blur (as shown in the zoomed-in area of Fig. 8(a)). These results are characterized by indistinct boundaries between tissues and the presence of noticeable noise, which compromises the clarity and distinctness of the reconstructed images. In contrast, the reconstruction of DPER with Tweedie denoising is much clearer. The 90° LACT task shows a similar phenomenon; for example, in Fig. 8(b), the blue arrow indicates that Tweedie denoising provides more precise information. The DPER that incorporates Tweedie demonstrates superior performance over the DPER without Tweedie for both the LACT and SVCT tasks. Recently, a lot of work has been demonstrating that using Tweedie denoising helps improve the performance of solving inverse problems via diffusion model [45, 65]. Indeed, using estimated $\mathbf{x}_{0}$ for data consistency operations (i.e., $\left\|\mathbf{A}\hat{\mathbf{x}}_{0|t}-\mathbf{y}\right\|_{2}^{2}$)) has become a widely used and recognized paradigm in the field of solving inverse problems via diffusion model [45, 65, 64, 61].
@@ 4.9 Effectiveness of Alternating Optimization Strategy
To evaluate the effectiveness of alternative optimization, we compared two naive strategies (as shown in Fig. 9) that directly integrate the INR and diffusion model [72, 73] without alternative optimization of the proposed framework. Strategy A: INR is first parameterized by directly fitting the incomplete measurement data. Then, the pre-trained diffusion model performs conditional sampling to refine the image provided by INR, producing the reconstruction result. Strategy B: First, the pre-trained diffusion model performs conditional sampling based on the FBP result of the incomplete measurement. Then, the generated result is refined by INR, followed by the Implicit Neural Representation Refinement process of DPER, to produce the reconstruction result. The conditional sampling process in Strategy B is conducted by adding an appropriate level of noise to the input FBP image, followed by reverse diffusion sampling.
@@ 4.10.2 Influence of Noises
During the process of CT acquisition, the presence of noise is an inevitable phenomenon, attributable to a myriad of factors. We conducted noise interference experiments by introducing noise to the projection data to verify the robustness of the proposed method. Specifically, we performed experiments on the LACT task with a 90° range and the SVCT task with 20 views, introducing Poisson noise (modeling photon variability noise) and Gaussian noise (modeling electronic system readout noise), respectively. For Poisson noise, we accomplished this through the use of a statistical Poisson model, as delineated below:
@@
$\mathbf{Y(\theta,\rho)}$ represents the transmitted X-ray photon intensity, where $b$ denotes the incident X-ray photon intensity, and $r$ symbolizes the average of the background events and read-out noise variance. To simulate different Signal-to-Noise Ratio (SNR) levels, we assign the value of $r$ as 10, while $b$ is set at $1.3\times 10^{4}$, $4\times 10^{4}$, and $4\times 10^{5}$, corresponding to approximate SNR levels of 32dB, 37 dB, and 43dB, respectively. The resulting noisy sinograms, denoted as $\mathbf{y}_{\theta}^{\prime}$, are computed using the negative logarithm transformation: $\mathbf{y}_{\theta}^{\prime}=-\ln(\mathbf{Y}(\theta,\rho)/b)$. For Gaussian noise, we followed the setting in [50], adding Gaussian noise with zero mean and 0.64 variance to the projected data. In our study, we evaluate the performance of our proposed DPER method against three baseline methodologies: FBP, SCOPE (as the SOTA in INR-based methods), and DiffusionMBIR (as the SOTA in diffusion-based methods) for CT image recovery from these noisy sinograms.
Figure 11: Qualitative and quantitative results of different methods to compare the measurement with Poisson noise ($b=1.3\times 10^{4}$) and Gaussian noise ($\sigma^{2}=0.64$) on one test sample (#18) of the AAPM dataset for the LACT task of [0, 90]° scanning range and SVCT task of 20 views.
Fig. 11 and Table 4 display the quantitative and qualitative results of the impact of different noise levels and types on our framework and other baselines for observing the robustness. There are some observations: 1) The results indicate that noise affects the model’s performance as it leads to a more ill-posed LACT inverse imaging problem. It is evident that as the noise level increases, the performance of all models deteriorates. However, the reduction ratio of data-specific models shows better results than Prior-based methods, which is expected (The effect of diffusion prior is limited by noise measurement). 2) Compared to the FBP result with noise, the DPER result is significantly cleaner and maintains the data fidelity, which represents that the DPER framework can adapt to a single instance with different noise levels and types. 3) Compared to the SOTA DiffusionMBIR, our framework maintains an improvement of nearly +2 dB PSNR in all noise settings for the LACT task and +1.5 dB PSNR for the SVCT task. The robustness of our method to noise is further demonstrated by the results in Fig. 11. The DiffusionMBIR results exhibit significant noise, resulting in blurred edges. In contrast, DPER stands out through effective noise reduction and clear structural details.
@@
Table 4: Quantitative results (PSNR / SSIM) of different methods on the AAPM dataset test samples for the LACT task with a [0, 90]° scanning range and the SVCT task with 20 views under different noise conditions (Poisson noise with SNR = 32 dB, 37 dB, 43 dB, and Gaussian noise with $\sigma^{2}=0.64$). The best performances are highlighted in bold.
Task Method Poisson Noise Gaussian Noise
32 dB 37 dB 43 dB $\sigma^{2}=0.64$
LACT [0,90]° FBP[67] 11.43 / 0.2018 11.47 / 0.2366 11.51 / 0.2974 12.36 / 0.2565
SCOPE[31] 22.75 / 0.6829 23.08 / 0.7409 23.32 / 0.8124 21.38 / 0.6867
DiffusionMBIR[46] 31.88 / 0.9234 33.07 / 0.9390 34.15 / 0.9503 31.65 / 0.9210
DPER (Ours) 33.36 / 0.9459 34.52 / 0.9602 36.41 / 0.9685 33.29 / 0.9584
SVCT 20 views FBP[67] 15.82 / 0.1961 17.62 / 0.2711 20.48 / 0.4473 17.12 / 0.2367
SCOPE[31] 27.73 / 0.7758 28.31 / 0.8015 28.24 / 0.8152 27.81 / 0.7912
DiffusionMBIR[46] 35.01 / 0.8915 36.92 / 0.9280 38.97 / 0.9497 36.01 / 0.9072
DPER (Ours) 36.75 / 0.9517 37.99 / 0.9567 39.77 / 0.9661 37.45 / 0.9537